An item produced by a company is susceptible to two types

of defects namely, A and B. The probability that an item has defect A is 1/6. The probability that it has defect

B is 1/8 . Calculate the probability that the item has 

  1. (i) Both A and B defects
  2. (ii) Only defect A
  3. (iii) Only defect B
  4. (iv) One defect only
  5. (v) At least one defect
1
Expert's answer
2021-07-21T15:02:46-0400

1.(i) "P=\frac{1}{6}*\frac{1}{8}=\frac{1}{48}\approx0.0208."


2.(ii) "P=\frac{1}{6}*(1-\frac{1}{8})=\frac{7}{48}\approx 0.1458."


3(iii) "P=\frac{1}{8}*(1-\frac{1}{6})=\frac{5}{48}\approx0.1042."


4(iv) "P=\frac{7}{48}+\frac{5}{48}=\frac{12}{48}=\frac{1}{4}=0.25."


5.(v) "P=1-\frac{5}{6}*\frac{7}{8}=1-\frac{35}{48}=\frac{13}{48}\approx0.2708."